OpenStudy (anonymous):

simplify by removing factors of 1 5t^2+5t/15t^2+25t

OpenStudy (anonymous):

Well, there's a t in everything....

OpenStudy (anonymous):

sorry thats the way the problem is written

OpenStudy (anonymous):

So u could take that out....

OpenStudy (anonymous):

When it would seem that the top is a third of the bottom.

OpenStudy (anonymous):

Ooops, not right, going blind....

OpenStudy (anonymous):

Better take 5 out of everything as well...

OpenStudy (anonymous):

Then ur done.

OpenStudy (anonymous):

would it be -1

OpenStudy (anonymous):

or 3/5

OpenStudy (anonymous):

5t^2+5t/15t^2+25t take out t's 5t+5/15t+25 take out 5's t+1/3t+5

OpenStudy (id21):

\[\frac{5t^2+5t}{15t^2+25t}=\frac{5t(t+1)}{5t(3t+5)}=\frac{t+1}{3t+5}\]